# Cost-Aware Contrastive Routing for LLMs

### v2

@@ A.3 Routing as Recommendation
Routing can also be framed as a recommendation problem, wherein the input query plays the role of a "user," the pool of LLMs corresponds to "items," and past performance metrics form the implicit interaction history [100, 93]. However, unlike conventional recommender systems, routing has limited "user" features (i.e., input metadata), making label collection and generalization especially challenging [61, 49].
Matrix factorization, attention-based models, and graph neural networks are used in both recommenders and routers [62, 105, 25], reinforcing the close link between the two domains.
@@ A.4 Scaling Laws and Architecture Trends
Scaling laws [42, 47] describe predictable trends between model size, data, and performance, guiding the development of efficient LLM architectures. These insights have been extended to MoEs [24, 13], sparse models [26], and hybrid systems [28, 63], offering context for when routing or merging approaches might be most beneficial.
@@ A.5 Routing within MoE and Hybrid Architectures
Routing is a central mechanism in MoE models [36, 41, 75], where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer [24] and Mixtral [38] employ sparse activation to minimize cost.
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Approaches like UltraFuser [20], Branch-Train-MiX [81], and token-level fusion highlight recent advances in combining model specialization and flexibility.
@@ B.1 Notation and Preliminaries
Let the Bayes‑optimal router be $r^{\star}(x)=\mathop{\rm arg\,min}_{m\in[M]}\bigl[\gamma(x,h_{m})+\lambda c_{m}\bigr]$ and define the excess cost–error gap
@@
Hence $\Delta_{r^{\star}(x)}(x)=0$ and $\Delta_{m}(x)\geq 0$. For a query $x$ let $r_{k}(x)$ be the radius of the ball $B(x,r)\subset\mathbb{S}^{d-1}$ (in the cosine metric) that contains the $k$-th nearest training neighbor. If the marginal on $q(\mathcal{X})$ has a density, $\operatorname*{\mathbb{E}}\!\bigl[r_{k}(x)^{d}\bigr]\leq C_{1}k/n$ [15, 79].
@@
Let the empirical cost‑adjusted risk be
@@
For convenience write $S_{im}=q_{i}^{\top}e_{m}/\tau_{k}$ when $m\in\mathcal{B}_{k}$. The single‑query loss (Equation (8)) is
@@
Taking the gradient w.r.t. $q_{i}$ and setting it to zero gives $\sum_{m}(p_{im}-q_{im})e_{m}=0$. Project onto $q_{i}$: $\sum_{m}(p_{im}-q_{im})S_{im}=0.$ Fix $m^{+}\in\mathcal{P}_{ik}$, $m^{-}\in\mathcal{N}(i)$ with $c_{m^{+}}\leq c_{m^{-}}$. Because $p_{im^{-}}=0$ while $p_{im^{+}}>0$, the equality forces $q_{im^{+}}>q_{im^{-}}$, hence $S_{im^{+}}-\gamma c_{m^{+}}>S_{im^{-}}-\gamma c_{m^{-}}$. Re‑arranging yields $q_{i}^{\top}e_{m^{+}}>q_{i}^{\top}e_{m^{-}}$, establishing directional alignment.
@@ C.1 Logit-Footprint Descriptors
These tokens aren’t used for their meaning. They’re probes of each model’s output behavior. Even common words get scored differently across models (temperature, punctuation/number handling, style). By averaging over many prompts and steps, the descriptor captures overall model behavior, not any single word’s semantics. Also, a recent work [82] shows that LLMs exhibit stable, word-level idiosyncrasies (as the authors call them) that enable near-perfect model attribution using only the first few generated tokens (even after paraphrasing or translation), implying that common tokens still provide discriminative signals about a model’s predictive calibration.
@@ C.2 The Step from Equation 3 to Equation 5
Because $p_{S}$ is fixed for all experts, $H(p_{h},p_{S})$ is a stable, model-specific quantity that makes descriptors comparable across experts (“same yardstick”). It is not the true NLL under $p_{h}$, but it preserves differences between experts via $H(p_{h})$ and their mismatch to $p_{S}$ via $\mathrm{KL}(p_{h}\|p_{S})$. In practice we use deterministic (greedy) decoding to reduce variance; averaging over many prompts/tokens makes the empirical $\tilde{\ell}_{h}(x)$ closely track the expectation above. Figure 2 empirically validates this argument.
@@ C.3 Band-Specific Temperatures and Smoother Gradients
This is the sense in which band-specific temperatures yield smoother gradients; it is especially helpful in high-cost bands that otherwise have few positives and highly variable similarities. Without band-specific temperatures, the router can exhibit oscillatory updates on hard prompts (rare positives dominate, then vanish), slowing convergence and encouraging over-use of cheap experts. Empirically, we observed that bands and band-specific temperatures are important (Table 7)
@@ C.4 Dense Human Annotations
  • Positive sets $P(i)$ with preference structure. Replace binary "correct expert" labels with pairwise preferences (cheap‐and‐good $\succ$ expensive‐and‐similar $\succ$ clearly wrong), yielding band-aware positives and margin constraints. This can be implemented by expanding $P(i)$ and adding a lightweight pairwise ranking (DPO-style) regularizer within each cost band.
@@ D.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines designed to capture key routing strategies and their trade-offs. Specifically, we include UMR [40], a recent state-of-the-art technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling [48, 2], which frames routing as a bandit exploration-exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing [34], a strategy that selects models by explicitly considering the cost-accuracy Pareto frontier; and two extreme baselines—Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle (Clairvoyant Upper-Bound [40]), which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric gating methods (Parametric Softmax Router) inspired by classical mixture-of-experts architectures [62] and SoftMoE, which models router decisions via differentiable soft gating functions [64]. Collectively, these baselines enable us to rigorously assess whether our contrastive routing approach delivers meaningful improvements in performance, cost-efficiency, and generalization capabilities relative to existing strategies.
@@
We train our router and evaluate it on three datasets: EmbedLLM [105], MixInstruct [39], and RouterBench [34]. For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)), resulting in a 256-dimensional vector per model. For RouterBench, we sample 192 probes from its training set, ensuring these probes are excluded from the training data used for the contrastive router. We compute perplexity-based descriptors on RouterBench and use GPT-2 [66]. On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric, meaning an LLM response is classified strictly as correct or incorrect. For MixInstruct, we employ exponentiated BARTScore [97] as the evaluation metric, following the approach in prior work [40, 39].
We evaluate each routing strategy using a deferral curve [40] which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (refer to Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, serving as a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following [40] we employ evaluation metrics including Area Under the Deferral Curve (AUDC), Query-Normalized Cost (QNC), and peak accuracy. QNC is the minimum relative cost required to match the performance of the most accurate tested LLM.
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2[70] model as the embedding backbone ($\Phi(x)$ in Section3) across all experiments. Our trainable router component is a two-layer MLP, denoted as $g_{\theta}(.)$, which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset, excluding the probe examples from RouterBench. Training is performed for 10 epochs using the AdamW optimizer with a batch size of 512 and a learning rate of $5\times 10^{-4}$. For the cost spectrum loss (Equation (8)), we set the number of cost bands to $K=5$ and the negative cost penalty to $\lambda=0.1$. The hyperparameters for the linear schedule of band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. All training and descriptor extraction are done on RTX6000Ada GPUs with $\sim$48GB GPU memory.
@@ D.3.1 Ablation of Number of Cost bands
Figure 7: Deferral‑curve metrics across different numbers of cost bands.
@@
Table 8: Deferral‑curve metrics across different numbers of cost bands.
@@ D.3.2 Ablation of the Number of Neighbors
Increasing $k$ (the number of ANN neighbors selected before cost-aware scoring) generally provides modest improvements before reaching a plateau. (The baseline UMR [40] also included an ablation on $k$ in a K-NN router.) Beyond a small $K$, the gains in accuracy or AUDC become minimal, while both latency and the likelihood of choosing unnecessarily expensive experts increase. We selected a default of $K=4$ and it worked reasonably well so we did not do further hyperparameter tuning. We performed an ablation on a subset of embedllm prompts. The results are shown in Table 9 which are consistent with the trend observed in UMR.
@@ D.3.4 Ablation of Band-Specific Temperature Slope
Table 11: Ablation over the band slope $\alpha$ in $\tau_{k}=\tau_{\min}+\alpha\,\bar{c}_{k}$. Moderate $\alpha$ (0.1–0.25) maximizes AUDC and lowers the cost required to reach peak accuracy; large $\alpha$ oversmooths high-cost bands and reduces peak accuracy.
See Table 11 for an ablation of this hyperparameter. As $\alpha$ increases from 0 to a moderate value (0.10.10.250.25), AUDC improves and the cost@max_acc drops, indicating we reach peak quality at lower cost, while max_acc remains comparable. For larger $\alpha$ ($\geq 0.4$), the high-cost bands become oversmoothed, weakening discrimination among expensive experts; max_acc and AUDC decline despite further cost reductions. This validates our choice of adopting a moderate setting (default $\alpha{=}0.25$, with $0.1$ performing slightly better).
@@ D.3.5 Ablation of Cheapest Cost Band
$\tau_{\min}$ is the softmax temperature for the cheapest cost band in Equation 7; all other bands inherit $\tau_{k}=\tau_{\min}+\alpha\,\bar{c}_{k}$. A very small $\tau_{\min}$ makes the cheap-band softmax sharp (highly discriminative but prone to noisy, peaky gradients) while a larger $\tau_{\min}$ smooths the distribution, lowering variance but also blurring differences among cheap experts.
@@
Table 12: Ablation over the base temperature $\tau_{\min}$ (with $\alpha=0.25$). Moderate values improve AUDC and lower the cost required for peak or near-peak accuracy; very small or large values underperform.
@@ D.6 Statistical Significance
We perform full evaluation on Embedllm [105] using paired, prompt-level significance tests to concretely assess statistical significance. Specifically, we computed a paired bootstrap [22] (sampling 3,000 prompts with replacement, 5,000 times) to obtain a 95% confidence interval for $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (UMR is the best baseline overall). We also report a one-sided p-value for the hypothesis $\Delta>0$. Additionally, we ran McNemar’s [57] test at a matched budget (using the median of the combined cost grids) to compare per-prompt wins and losses at equivalent operating cost. These tests quantify uncertainty over the test prompts.
@@
Table 16: Paired significance versus the strongest baseline. $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (area under the deferral curve; higher is better). “95% CI” and the one-sided $p$ come from a paired bootstrap over prompts ($N=3000$, $B=5000$; $H_{1}\!:\Delta>0$). $c^{\star}$ is the matched budget used for McNemar. $n_{10}/n_{01}$ are discordant counts (CSCR correct / baseline correct), and “$p$ (McNemar)” is the one-sided exact binomial $p$ for CSCR >> baseline at $c^{\star}$. CIs that exclude 0 and small $p$-values indicate a statistically significant improvement of CSCR.

### v3

@@ A.3 Routing as Recommendation
Routing can also be framed as a recommendation problem, wherein the input query plays the role of a "user", the pool of LLMs corresponds to "items", and past performance metrics form the implicit interaction history [zhao2024recommender, wu2022disentangled]. However, unlike conventional recommender systems, routing has limited "user" features (i.e., input metadata), making label collection and generalization especially challenging [nguyen2024metallm, liu2024optllm].
Matrix factorization, attention-based models, and graph neural networks are used in both recommenders and routers [OngAlmWu2024, zhuang2024embedllm, Feng:2024], reinforcing the close link between the two domains.
@@ A.4 Scaling Laws and Architecture Trends
Scaling laws [kaplan2020scaling, li2025mis] describe predictable trends between model size, data, and performance, guiding the development of efficient LLM architectures. These insights have been extended to MoEs [fedus2022switch, clark2022unified], sparse models [frantar2023scaling], and hybrid systems [gu2023mamba, poli2024mechanistic], offering context for when routing or merging approaches might be most beneficial.
@@ A.5 Routing within MoE and Hybrid Architectures
Routing is a central mechanism in MoE models [Jacobs:1991, Jordan:1993, outrageously], where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer [fedus2022switch] and Mixtral [jiang2024mixtral] employ sparse activation to minimize cost.
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Approaches like UltraFuser [ding2024mastering], Branch-Train-MiX [sukhbaatar2024branch], and token-level fusion highlight recent advances in combining model specialization and flexibility.
@@ B.1 Notation and Preliminaries
Let the Bayes-optimal router be $r^{\star}(x)=\mathop{\rm arg\,min}_{m\in[M]}\bigl[\gamma(x,h_{m})+\lambda c_{m}\bigr]$ and define the excess cost–error gap
@@
Hence $\Delta_{r^{\star}(x)}(x)=0$ and $\Delta_{m}(x)\geq 0$. For a query $x$ let $r_{k}(x)$ be the radius of the ball $B(x,r)\subset\mathbb{S}^{d-1}$ (in the cosine metric) that contains the $k$-th nearest training neighbor. If the marginal on $q(\mathcal{X})$ has a density, $\operatorname*{\mathbb{E}}\!\bigl[r_{k}(x)^{d}\bigr]\leq C_{1}k/n$ [cover1967nearest, stone1977consistent].
@@
Let the empirical cost-adjusted risk be
@@
For convenience write $S_{im}=q_{i}^{\top}e_{m}/\tau_{k}$ when $m\in\mathcal{B}_{k}$. The single-query loss (Equation (8)) is
@@
Taking the gradient w.r.t. $q_{i}$ and setting it to zero gives $\sum_{m}(p_{im}-q_{im})e_{m}=0$. Project onto $q_{i}$: $\sum_{m}(p_{im}-q_{im})S_{im}=0.$ Fix $m^{+}\in\mathcal{P}_{ik}$, $m^{-}\in\mathcal{N}(i)$ with $c_{m^{+}}\leq c_{m^{-}}$. Because $p_{im^{-}}=0$ while $p_{im^{+}}>0$, the equality forces $q_{im^{+}}>q_{im^{-}}$, hence $S_{im^{+}}-\gamma c_{m^{+}}>S_{im^{-}}-\gamma c_{m^{-}}$. Rearranging yields $q_{i}^{\top}e_{m^{+}}>q_{i}^{\top}e_{m^{-}}$, establishing directional alignment.
@@ C.1 Logit-Footprint Descriptors
These tokens aren’t used for their meaning. They’re probes of each model’s output behavior. Even common words get scored differently across models (temperature, punctuation/number handling, style). By averaging over many prompts and steps, the descriptor captures overall model behavior, not any single word’s semantics. Also, a recent work [sun2025idiosyncrasies] shows that LLMs exhibit stable, word-level idiosyncrasies (as the authors call them) that enable near-perfect model attribution using only the first few generated tokens (even after paraphrasing or translation), implying that common tokens still provide discriminative signals about a model’s predictive calibration.
@@ C.2 The Step from Equation 3 to Equation 5
Because $p_{S}$ is fixed for all experts, $H(p_{h},p_{S})$ is a stable, model-specific quantity that makes descriptors comparable across experts (“same yardstick”). It is not the true NLL under $p_{h}$, but it preserves differences between experts via $H(p_{h})$ and their mismatch to $p_{S}$ via $\mathrm{KL}(p_{h}\|p_{S})$. In practice we use deterministic (greedy) decoding to reduce variance, averaging over many prompts/tokens makes the empirical $\tilde{\ell}_{h}(x)$ closely track the expectation above. Figure 2 empirically validates this argument.
@@ C.3 Band-Specific Temperatures and Smoother Gradients
This is the sense in which band-specific temperatures yield smoother gradients. It is especially helpful in high-cost bands that otherwise have few positives and highly variable similarities. Without band-specific temperatures, the router can exhibit oscillatory updates on hard prompts (rare positives dominate, then vanish), slowing convergence and encouraging over-use of cheap experts. Empirically, we observed that bands and band-specific temperatures are important (Table 7)
@@ C.4 Dense Human Annotations
  • Positive sets $P(i)$ with preference structure. Replace binary "correct expert" labels with pairwise preferences (cheap-and-good $\succ$ expensive-and-similar $\succ$ clearly wrong), yielding band-aware positives and margin constraints. This can be implemented by expanding $P(i)$ and adding a lightweight pairwise ranking (DPO-style) regularizer within each cost band.
@@ D.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines designed to capture key routing strategies and their trade-offs. Specifically, we include UMR [jitkrittum2025umr], a recent state-of-the-art technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling [Li:2025, agrawal2012analysis-thoompson], which frames routing as a bandit exploration–exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing [HuBieLi2024], a strategy that selects models by explicitly considering the cost–accuracy Pareto frontier; and two extreme baselines—Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle (Clairvoyant Upper-Bound [jitkrittum2025umr]), which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric gating methods (Parametric Softmax Router) inspired by classical mixture-of-experts architectures [OngAlmWu2024] and SoftMoE, which models router decisions via differentiable soft gating functions [puigcerver2024soft-moe]. Collectively, these baselines enable us to rigorously assess whether our contrastive routing approach delivers meaningful improvements in performance, cost-efficiency, and generalization capabilities relative to existing strategies.
@@
We train our router and evaluate it on three datasets: EmbedLLM [zhuang2024embedllm], MixInstruct [llmblender], and RouterBench [HuBieLi2024]. For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)), resulting in a 256-dimensional vector per model. For RouterBench, we sample 192 probes from its training set, ensuring these probes are excluded from the training data used for the contrastive router. We compute perplexity-based descriptors on RouterBench and use GPT-2 [radford2019language]. On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric, meaning an LLM response is classified strictly as correct or incorrect. For MixInstruct, we employ exponentiated BARTScore [yuan2021bartscore] as the evaluation metric, following the approach in prior work [jitkrittum2025umr, llmblender].
We evaluate each routing strategy using a deferral curve [jitkrittum2025umr] which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (refer to Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, serving as a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following [jitkrittum2025umr] we employ evaluation metrics including Area Under the Deferral Curve (AUDC), Query-Normalized Cost (QNC), and peak accuracy. QNC is the minimum relative cost required to match the performance of the most accurate tested LLM.
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2[reimers-2019-sentence-bert] model as the embedding backbone ($\Phi(x)$ in Section3) across all experiments. Our trainable router component is a two-layer MLP, denoted as $g_{\theta}(.)$, which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset, excluding the probe examples from RouterBench. Training is performed for 10 epochs using the AdamW optimizer with a batch size of 512 and a learning rate of $5\times 10^{-4}$. For the cost spectrum loss (Equation (8)), we set the number of cost bands to $K=5$ and the negative cost penalty to $\lambda=0.1$. The hyperparameters for the linear schedule of band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. All training and descriptor extraction are done on RTX6000Ada GPUs with $\sim$48GB GPU memory.
@@ D.3.1 Ablation of Number of Cost bands
Figure 7: Deferral-curve metrics across different numbers of cost bands.
@@
Table 8: Deferral-curve metrics across different numbers of cost bands.
@@ D.3.2 Ablation of the Number of Neighbors
Increasing $k$ (the number of ANN neighbors selected before cost-aware scoring) generally provides modest improvements before reaching a plateau. (The baseline UMR [jitkrittum2025umr] also included an ablation on $k$ in a K-NN router.) Beyond a small $K$, the gains in accuracy or AUDC become minimal, while both latency and the likelihood of choosing unnecessarily expensive experts increase. We selected a default of $K=4$ and it worked reasonably well so we did not do further hyperparameter tuning. We performed an ablation on a subset of embedllm prompts. The results are shown in Table 9 which are consistent with the trend observed in UMR.
@@ D.3.4 Ablation of Band-Specific Temperature Slope
Table 11: Ablation over the band slope $\alpha$ in $\tau_{k}=\tau_{\min}+\alpha\,\bar{c}_{k}$. Moderate $\alpha$ (0.1–0.25) maximizes AUDC and lowers the cost required to reach peak accuracy. Large $\alpha$ oversmooths high-cost bands and reduces peak accuracy.
See Table 11 for an ablation of this hyperparameter. As $\alpha$ increases from 0 to a moderate value (0.10.10.250.25), AUDC improves and the cost@max_acc drops, indicating we reach peak quality at lower cost, while max_acc remains comparable. For larger $\alpha$ ($\geq 0.4$), the high-cost bands become oversmoothed, weakening discrimination among expensive experts so max_acc and AUDC decline despite further cost reductions. This validates our choice of adopting a moderate setting (default $\alpha{=}0.25$, with $0.1$ performing slightly better).
@@ D.3.5 Ablation of Cheapest Cost Band
$\tau_{\min}$ is the softmax temperature for the cheapest cost band in Equation 7. All other bands inherit $\tau_{k}=\tau_{\min}+\alpha\,\bar{c}_{k}$. A very small $\tau_{\min}$ makes the cheap-band softmax sharp (highly discriminative but prone to noisy, peaky gradients) while a larger $\tau_{\min}$ smooths the distribution, lowering variance but also blurring differences among cheap experts.
@@
Table 12: Ablation over the base temperature $\tau_{\min}$ (with $\alpha=0.25$). Moderate values improve AUDC and lower the cost required for peak or near-peak accuracy. Very small or large values underperform.
@@ D.6 Statistical Significance
We perform full evaluation on Embedllm [zhuang2024embedllm] using paired, prompt-level significance tests to concretely assess statistical significance. Specifically, we computed a paired bootstrap [efron1994bootstrap] (sampling 3,000 prompts with replacement, 5,000 times) to obtain a 95% confidence interval for $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (UMR is the best baseline overall). We also report a one-sided p-value for the hypothesis $\Delta>0$. Additionally, we ran McNemar’s [mcnemar1947note] test at a matched budget (using the median of the combined cost grids) to compare per-prompt wins and losses at equivalent operating cost. These tests quantify uncertainty over the test prompts.
@@
Table 16: Paired significance versus the strongest baseline. $\Delta_{\mathrm{AUDC}}=\mathrm{AUDC}_{\text{CSCR}}-\mathrm{AUDC}_{\text{UMR}}$ (area under the deferral curve. Higher is better). “95% CI” and the one-sided $p$ come from a paired bootstrap over prompts ($N=3000$, $B=5000$, $H_{1}\!:\Delta>0$). $c^{\star}$ is the matched budget used for McNemar. $n_{10}/n_{01}$ are discordant counts (CSCR correct / baseline correct), and “$p$ (McNemar)” is the one-sided exact binomial $p$ for CSCR >> baseline at $c^{\star}$. CIs that exclude 0 and small $p$-values indicate a statistically significant improvement of CSCR.